The lines described by y=(4a-6)x and y=1/5x are perpendicular. What is the value of a?
Answer by macston(5194)

.
The slope of the line y=(1/5)x is 1/5.
The slope of the perpendicular line is the negative reciprocal=-(5/1)=-5
For y=(4a-6)x, (4a-6) is the slope, so to find a:
.
4a-6=-5 Add 6 to each side.
4a=1 Divide each side by 4.
a=1/4
ANSWER: For the line perpendicular to y=(1/5)x, a=1/4.